Let y = l^2-l^3/z where l = 2.0 + 0.1, z = 1.0 + 0.1 then the value of y is given by
    (a)    + 2 + 0.8
    (b)    –4 + 1.6
    (c)    –4 + 0.8
    (d)    + 2 + 1.6

Grade:10

1 Answers

Navjyot Kalra
10 years ago

(b)

    y = l^2-l^3/z

    dy = 2ldl-((z.3l^2 dl-l^3 dz)/z^2 )=(2l-(3l^2)/z)dl+l^3/z^2  dz

    = (2×2-(3×2^2)/1)(±0.1)+8/1 (±0.1)

    = – 8 (+0.1) +8 (+0.01) = + 1.6

    y = l^2-l^3/z=2^2-2^3/1 = 4 – 8 = –4

    => y = – 4 + 1.6